# Connected Components in Linear Work and Near-Optimal Time

### v2

@@ 4.1 A Constant-Shrink Algorithm
The algorithm is inspired by the constant-shrink MPC algorithm from [${}^{+}$19], which uses the $O(1)$-time sorting and prefix sum computing on an MPC. We could also modify Gazit’s random-mate algorithm [Gaz91] to get a constant-shrink algorithm. We include our version here tuned for our applications and for completeness.
Matching$(E)$: 1. For each edge $(u,v)\in E$: if $u$ or $v$ is not a root or $u=v$ then delete $(u,v)$ from $E$. 2. For each edge $e\in E$: orient it from the large end to the small end. This creates a digraph $D$. 3. For each vertex $v\in V(E)$: if $v$ has more than $1$ outgoing arcs then keep an arbitrary one and delete the others. 4. For each singleton $v$ created in Step 3: consider $D$ before Step 3 and choose an arbitrary arc $(u,v)$ in $D$ then $v.p=u$. 5. For each root $v\in V(E)$: if $v$ has more than $1$ incoming arcs then remove all outgoing arcs of $v$. 6. For each root $v\in V(E)$: if $v$ has more than $1$ incoming arcs then update the parents of all vertices with arcs to $v$ as $v$ and delete all these vertices from $D$. 7. For each arc $e$ in $D$: w.p. $1/2$ delete $e$ from $D$. 8. For each arc $(u,v)$ in $D$: if $(u,v)$ is isolated (not sharing ends with other arcs), then $v.p=u$. 9. For each vertex $v\in V(E)$: $v.p=v.p.p$.
@@ 4.2 A $\log\log n$-Shrink Algorithm
Alter$(E)$: 1. For each edge $(u,v)\in E$: replace $(u,v)$ by $(u.p,v.p)$ in $E$. 2. Remove loops from $E$.
@@
Filter$(E,k)$: 1. For round $j$ from 00 to $k$: Matching$(E)$, Alter$(E)$, delete each edge from $E$ w.p. $10^{-4}$. 2. For iteration $j$ from $k$ to 00: if a vertex $v$ updates $v.p$ in round $j$ then $v.p=v.p.p$. 3. Return $V(E)$.
@@
Reverse$(V^{\prime},E)$: 1. For each non-root vertex $v\in V^{\prime}$: $v.p.p=v$, $v.p=v.p.p$. 2. For each vertex $v$ in the labeled digraph: $v.p=v.p.p$. 3. Alter$(E)$.
@@ 4.3 A $\operatorname{poly}(\log n)$-Shrink Algorithm
Reduce$(V,E,k)$: 1. Extract$(E,1000\log\log\log n)$. 2. $V^{\prime}=\textsc{Filter}(E,k)$. 3. For each $v\in V$ do $v.p=v.p.p$. Alter$(E)$. 4. Edge set $E^{\prime}=\{(u,v)\mid u\notin V^{\prime}\text{ or }v\notin V^{\prime}\}$. 5. For round $i$ from 00 to $k$: Matching$(E^{\prime})$, for each $v\in V$ do $v.p=v.p.p$, Alter$(E^{\prime})$. 6. Reverse$(V^{\prime},E)$.
@@ 5 Stage 2: Increasing the Minimum Degree to $\operatorname{poly}(\log n)$
Increase$(V,E,b)$: 1. Skeleton graph $H=(V,E^{\prime})=\textsc{Build}(V,E,b)$. 2. $(V,E_{\text{close}})=\textsc{Densify}(H,b)$. 3. Use $E_{\text{close}}$ and the labeled digraph updated in Step 2 to increase the degree of each root (vertex in the current graph) to be at least $\text{poly}(b)$ – see §5.3 for a detailed description.
@@ 5.1 Building the Skeleton Graph
Build$(V,E,b)$: 1. For each vertex $v\in V$: assign a block of $b^{9}$ processors, which will be used as an indexed hash table $\mathcal{H}(v)$ of size $b^{9}$. 2. Choose a random hash function $h:[n]\to[b^{9}]$. For each edge $(u,v)\in E$: write $u$ into the $h(u)$-th cell of $\mathcal{H}(v)$. 3. For each vertex $v\in V$: if the number of vertices in $\mathcal{H}(v)$ is more than $b^{8}$ then mark $v$ as high, else mark $v$ as low. 4. Let $E^{\prime}$ be an empty edge set. For each edge $(u,v)\in E$: if $u$ or $v$ is low then add $(u,v)$ to $E^{\prime}$, else add $(u,v)$ to $E^{\prime}$ w.p. $b^{-1}$. 5. Remove parallel edges and loops from $E^{\prime}$. 6. Return graph $H=(V,E^{\prime})$.
@@ 5.2.1 Algorithmic Framework of Densify
Expand-Maxlink$(H)$: 1. Let $V=V(H)$ and $E=E(H)$. 2. Maxlink$(V)$; Alter$(E)$. 3. For each root $v\in V$: increase $\ell(v)$ by $1$ w.p. $\beta(v)^{-0.06}$. 4. For each root $v\in V$: for each root $w\in N^{*}(v)$: if $\beta(w)=\beta(v)$ then hash $w$ into $\mathcal{H}(v)$. 5. For each root $v\in V$: if there is a collision in $\mathcal{H}(v)$ then mark $v$ as dormant. For each vertex $v\in V$: if there is a dormant vertex in $\mathcal{H}(v)$ then mark $v$ as dormant. 6. For each root $v\in V$: for each $w\in\mathcal{H}(v)$: for each $u\in\mathcal{H}(w)$: hash $u$ into $\mathcal{H}(v)$. For each root $v\in V$: if there is a collision in $\mathcal{H}(v)$ then mark $v$ as dormant. 7. Maxlink$(V)$; Shortcut$(V)$; Alter$(E(V))$. 8. For each root $v\in V$: if $v$ is dormant and did not increase its level in Step 3 then increase $\ell(v)$ by $1$. 9. For each root $v\in V$: assign a block of size $\beta_{\ell(v)}$ to $v$. 10. Return $H$ as the current graph with all added edges (in the form of items in all hash tables).
Densify$(H,b)$: 1. For round $i$ from 00 to $20\log b$ : $H=\textsc{Expand-Maxlink}(H)$. 2. Let $V$ be the set of roots in $H$. 3. Repeat for $2\log\log\log n$ times: Shortcut$(V(H))$. Alter$(E(H))$. 4. Let $E_{\text{close}}$ be the set of edges in the current graph $H$, including all (altered) edges from the input graph and all (altered) edges in all hash tables of all vertices. 5. Call the algorithm in Theorem 2 on graph $(V(E_{\text{close}}),E_{\text{close}})$ to update the labeled digraph. 6. Alter$(E_{\text{close}})$. 7. Return $(V,E_{\text{close}})$.
@@ 5.2.2 Properties of Densify
[…]
@@ 5.2.3 Path Construction for Densify
PathConstruction$(H,P)$: Let all vertices on $P$ be active at the beginning of round 00 of Densify$(H,b)$. For any round $i\in[0,9\log b)$, given path $P_{i}$ with at least $4$ active vertices at the beginning of round $i$, Expand-Maxlink$(H)$ constructs $P_{i+1}$ by the following $6$ phases: 1. The Alter$(E)$ in Step 2 replaces each vertex $v$ on $P_{i}$ by $v^{\prime}\coloneqq v.p$ to get path $P_{i,1}$. Set $v^{\prime}$ as active or passive according to $v$. For any $v^{\prime}$ on $P_{i,1}$, let $\underline{v^{\prime}}$ be on $P_{i}$ such that $\underline{v^{\prime}}.p=v^{\prime}$. 2. After Step 6, set $j$ as $1$, and repeat the following until $j\geq|P_{i,1}|-1$ or $P_{i,1}(j+1)$ is passive (not active): Let $v^{\prime}\coloneqq P_{i,1}(j)$, if $\underline{v^{\prime}}$ is a root and does not increase level during round $i$ then: (a) If the current graph contains edge $(v^{\prime},P_{i,1}(j+2))$ then mark $P_{i,1}(j+1)$ as skipped and set $j\coloneqq j+2$. (b) Else set $j\coloneqq j+1$. 3. Let $k$ be the maximum integer such that $P_{i,1}(k)$ is active and $k\leq j$. Mark $P_{i,1}(k)$ as passive. 4. Remove all skipped and passive vertices from $P_{i,1}$ to get path $P_{i,4}$. 5. Concatenate $P_{i,4}$ with all passive vertices after $P_{i,1}(|P_{i,4}|)$ (exclusively) on $P_{i,1}$ to get path $P_{i,5}$. 6. The Alter$(E)$ in Step 7 replaces each vertex $v$ on $P_{i,5}$ by $v.p$ to get path $P_{i+1}$. Set $v.p$ as active or passive according to $v$. For any vertex $v$ on $P_{i}$ that is replaced by $v^{\prime}$ in Phase 1, if $v^{\prime}$ is not removed in Phase 4, then let $\overline{v}$ be the vertex replacing $v^{\prime}$ in Phase 6, and call $\overline{v}$ the corresponding vertex of $v$ in round $i+1$.
@@
PassivePathConstruction$(H,P)$: Let $r=9\log b-1$. For any integer $i\in[0,r]$, let path $P_{i}$ be defined as in PathConstruction$(H,P)$. For any integer $i\in[r,20\log b]$, given path $P_{i}$ with at least $3$ vertices at the beginning of round $i$, Expand-Maxlink$(H)$ constructs $P_{i+1}$ by the following $3$ phases: 1. The Alter$(E)$ in Step 2 replaces each vertex $v$ on $P_{i}$ by $v^{\prime}\coloneqq v.p$ to get path $P_{i,1}$. For any $v^{\prime}$ on $P_{i,1}$, let $\underline{v^{\prime}}$ be on $P_{i}$ such that $\underline{v^{\prime}}.p=v^{\prime}$. 2. After Step 6, let $v^{\prime}\coloneqq P_{i,1}(0)$, i.e., the first vertex on $P_{i,1}$. If $\underline{v^{\prime}}$ is a root at the end of round $i$ and does not increase level during round $i$ then: if the current graph contains edge $(v^{\prime},P_{i,1}(2))$ then remove $P_{i,1}(1)$. After removing all such vertices from $P_{i,1}$ we obtain $P_{i,2}$. 3. The Alter$(E)$ in Step 7 replaces each vertex $v$ on $P_{i,2}$ by $v.p$ to get path $P_{i+1}$. For any vertex $v$ on $P_{i}$ that is replaced by $v^{\prime}$ in Phase 1, if $v^{\prime}$ is not removed in Phase 2, then let $\overline{v}$ be the vertex replacing $v^{\prime}$ in Phase 3, and call $\overline{v}$ the corresponding vertex of $v$ in round $i+1$.
@@
[…]
@@ 5.3.1 Algorithmic Framework of Increase
Increase$(V,E,b)$: 1. Subgraph $H=(V,E^{\prime})=\textsc{Build}(V,E,b)$. 2. $(V,E_{\text{close}})=\textsc{Densify}(H,b)$. 3. For each vertex $v\in V$: assign a block of $b^{9}$ processors, which will be used as an indexed hash table $\mathcal{H}^{\prime}(v)$ of size $b^{9}$. 4. Choose a random hash function $h:[n]\to[b^{9}]$. For each vertex $v\in V$: compute $u\coloneqq v.p^{(2R+1)}$, then write $v$ into the $h(v)$-th cell of $\mathcal{H}^{\prime}(u)$ and update $v.p=u$. 5. For each vertex $v\in V$: if the number of vertices in $\mathcal{H}^{\prime}(v)$ is at least $2b$ then mark $v$ as a head. 6. For each non-loop edge $(v,w)\in E_{\text{close}}$: if $v$ is a head and $w$ is a non-head then $w.p=v$. 7. Shortcut$(V)$. 8. For each vertex $v\in V$: sample $v$ as a leader w.p. $1/2$. For each non-loop edge $(v,w)\in E_{\text{close}}$: if $v$ is a leader and $w$ is a non-leader then $w.p.p=v.p$. 9. Shortcut$(V)$. 10. Alter$(E)$.

### v4

@@ 4.1 A Constant-Shrink Algorithm
The algorithm is inspired by the constant-shrink MPC algorithm from [BDE+19], which uses the $O(1)$-time sorting and prefix sum computing on an MPC. We could also modify Gazit’s random-mate algorithm [Gaz91] to get a constant-shrink algorithm. We include our version here tuned for our applications and for completeness.
Matching$(E)$:
  1. 1.
    For each edge $(u,v)\in E$: if $u$ or $v$ is not a root or $u=v$ then delete $(u,v)$ from $E$.
  2. 2.
    For each edge $e\in E$: orient it from the large end to the small end. This creates a digraph $D$.
  3. 3.
    For each vertex $v\in V(E)$: if $v$ has more than $1$ outgoing arcs then keep an arbitrary one and delete the others.
  4. 4.
    For each singleton $v$ created in Step 3: consider $D$ before Step 3 and choose an arbitrary arc $(u,v)$ in $D$ then $v.p=u$.
  5. 5.
    For each root $v\in V(E)$: if $v$ has more than $1$ incoming arcs then remove all outgoing arcs of $v$.
  6. 6.
    For each root $v\in V(E)$: if $v$ has more than $1$ incoming arcs then update the parents of all vertices with arcs to $v$ as $v$ and delete all these vertices from $D$.
  7. 7.
    For each arc $e$ in $D$: w.p. $1/2$ delete $e$ from $D$.
  8. 8.
    For each arc $(u,v)$ in $D$: if $(u,v)$ is isolated (not sharing ends with other arcs), then $v.p=u$.
  9. 9.
    For each vertex $v\in V(E)$: $v.p=v.p.p$.
@@ 4.2 A $\log\log n$-Shrink Algorithm
Alter$(E)$:
  1. 1.
    For each edge $(u,v)\in E$: replace $(u,v)$ by $(u.p,v.p)$ in $E$.
  2. 2.
    Remove loops from $E$.
@@
Filter$(E,k)$:
  1. 1.
    For round $j$ from 00 to $k$: Matching$(E)$, Alter$(E)$, delete each edge from $E$ w.p. $10^{-4}$.
  2. 2.
    For iteration $j$ from $k$ to 00: if a vertex $v$ updates $v.p$ in round $j$ then $v.p=v.p.p$.
  3. 3.
    Return $V(E)$.
@@
Reverse$(V^{\prime},E)$:
  1. 1.
    For each non-root vertex $v\in V^{\prime}$: $v.p.p=v$, $v.p=v.p.p$.
  2. 2.
    For each vertex $v$ in the labeled digraph: $v.p=v.p.p$.
  3. 3.
    Alter$(E)$.
@@
  2. 2.
    For round $i$ from 00 to $k$: $V^{\prime}=V^{\prime}+\textsc{Filter}(E^{\prime},k)$, Alter$(E^{\prime})$, $E^{\prime}=E^{\prime}-\{(u,v)\mid u,v\in V^{\prime}\}$.
  3. 3.
    For iteration $i$ from $k$ to 00: if a vertex $v$ updates $v.p$ in round $i$ then $v.p=v.p.p$.
  4. 4.
    Reverse$(V^{\prime},E)$.
@@ 4.3 A $\operatorname{poly}(\log n)$-Shrink Algorithm
Reduce$(V,E,k)$:
  1. 1.
    Extract$(E,1000\log\log\log n)$.
  2. 2.
    $V^{\prime}=\textsc{Filter}(E,k)$.
  3. 3.
    For each $v\in V$ do $v.p=v.p.p$. Alter$(E)$.
  4. 4.
    Edge set $E^{\prime}=\{(u,v)\mid u\notin V^{\prime}\text{ or }v\notin V^{\prime}\}$.
  5. 5.
    For round $i$ from 00 to $k$: Matching$(E^{\prime})$, for each $v\in V$ do $v.p=v.p.p$, Alter$(E^{\prime})$.
  6. 6.
    Reverse$(V^{\prime},E)$.
@@ 5 Stage 2: Increasing the Minimum Degree to $\operatorname{poly}(\log n)$
Increase$(V,E,b)$:
  1. 1.
    Skeleton graph $H=(V,E^{\prime})=\textsc{Build}(V,E,b)$.
  2. 2.
    $(V,E_{\text{close}})=\textsc{Densify}(H,b)$.
  3. 3.
    Use $E_{\text{close}}$ and the labeled digraph updated in Step 2 to increase the degree of each root (vertex in the current graph) to be at least $\text{poly}(b)$ – see §5.3 for a detailed description.
@@ 5.1 Building the Skeleton Graph
Build$(V,E,b)$:
  1. 1.
    For each vertex $v\in V$: assign a block of $b^{9}$ processors, which will be used as an indexed hash table $\mathcal{H}(v)$ of size $b^{9}$.
  2. 2.
    Choose a random hash function $h:[n]\to[b^{9}]$. For each edge $(u,v)\in E$: write $u$ into the $h(u)$-th cell of $\mathcal{H}(v)$.
  3. 3.
    For each vertex $v\in V$: if the number of vertices in $\mathcal{H}(v)$ is more than $b^{8}$ then mark $v$ as high, else mark $v$ as low.
  4. 4.
    Let $E^{\prime}$ be an empty edge set. For each edge $(u,v)\in E$: if $u$ or $v$ is low then add $(u,v)$ to $E^{\prime}$, else add $(u,v)$ to $E^{\prime}$ w.p. $b^{-1}$.
  5. 5.
    Remove parallel edges and loops from $E^{\prime}$.
  6. 6.
    Return graph $H=(V,E^{\prime})$.
@@ 5.2.1 Algorithmic Framework of Densify
Expand-Maxlink$(H)$:
  1. 1.
    Let $V=V(H)$ and $E=E(H)$.
  2. 2.
    Maxlink$(V)$; Alter$(E)$.
  3. 3.
    For each root $v\in V$: increase $\ell(v)$ by $1$ w.p. $\beta(v)^{-0.06}$.
  4. 4.
    For each root $v\in V$: for each root $w\in N^{*}(v)$: if $\beta(w)=\beta(v)$ then hash $w$ into $\mathcal{H}(v)$.
  5. 5.
    For each root $v\in V$: if there is a collision in $\mathcal{H}(v)$ then mark $v$ as dormant. For each vertex $v\in V$: if there is a dormant vertex in $\mathcal{H}(v)$ then mark $v$ as dormant.
  6. 6.
    For each root $v\in V$: for each $w\in\mathcal{H}(v)$: for each $u\in\mathcal{H}(w)$: hash $u$ into $\mathcal{H}(v)$. For each root $v\in V$: if there is a collision in $\mathcal{H}(v)$ then mark $v$ as dormant.
  7. 7.
    Maxlink$(V)$; Shortcut$(V)$; Alter$(E(V))$.
  8. 8.
    For each root $v\in V$: if $v$ is dormant and did not increase its level in Step 3 then increase $\ell(v)$ by $1$.
  9. 9.
    For each root $v\in V$: assign a block of size $\beta_{\ell(v)}$ to $v$.
  10. 10.
    Return $H$ as the current graph with all added edges (in the form of items in all hash tables).
Densify$(H,b)$:
  1. 1.
    For round $i$ from 00 to $20\log b$ : $H=\textsc{Expand-Maxlink}(H)$.
  2. 2.
    Let $V$ be the set of roots in $H$.
  3. 3.
    Repeat for $2\log\log\log n$ times: Shortcut$(V(H))$. Alter$(E(H))$.
  4. 4.
    Let $E_{\text{close}}$ be the set of edges in the current graph $H$, including all (altered) edges from the input graph and all (altered) edges in all hash tables of all vertices.
  5. 5.
    Call the algorithm in Theorem 2 on graph $(V(E_{\text{close}}),E_{\text{close}})$ to update the labeled digraph.
  6. 6.
    Alter$(E_{\text{close}})$.
  7. 7.
    Return $(V,E_{\text{close}})$.
@@ 5.2.2 Properties of Densify
W.p. $1-1/(\log n)^{8}$, at the end of Densify$(H,b)$, for any vertex $v\in V(H)$, $v$ is either a root or a child of a root, [^12] We did not state that all trees in the labeled digraph are flat because the vertex set $V(H)$ does not contain all the vertices in the original input graph: $H$ is the current graph after Stage $1$ that induced on roots. The non-roots created in Stage $1$ can be as many as $\Theta(n)$ so that we cannot run Shortcut on all of them. Instead, we will prove that at the end of Stage $3$, all such non-roots are children or grandchildren of roots so that one application of Shortcut flattens all trees and thus gives a correct connected components algorithm, which adds $O(n)$ work in total. and both ends of any edge in $E_{\text{close}}$ are roots.
@@ 5.2.3 Path Construction for Densify
PathConstruction$(H,P)$:
Let all vertices on $P$ be active at the beginning of round 00 of Densify$(H,b)$.
For any round $i\in[0,9\log b)$, given path $P_{i}$ with at least $4$ active vertices at the beginning of round $i$, Expand-Maxlink$(H)$ constructs $P_{i+1}$ by the following $6$ phases:
  1. 1.
    The Alter$(E)$ in Step 2 replaces each vertex $v$ on $P_{i}$ by $v^{\prime}\coloneqq v.p$ to get path $P_{i,1}$. Set $v^{\prime}$ as active or passive according to $v$. For any $v^{\prime}$ on $P_{i,1}$, let $\underline{v^{\prime}}$ be on $P_{i}$ such that $\underline{v^{\prime}}.p=v^{\prime}$.
  2. 2.
    After Step 6, set $j$ as $1$, and repeat the following until $j\geq|P_{i,1}|-1$ or $P_{i,1}(j+1)$ is passive (not active):
    • Let $v^{\prime}\coloneqq P_{i,1}(j)$, if $\underline{v^{\prime}}$ is a root and does not increase level during round $i$ then:
      1. (a)
        If the current graph contains edge $(v^{\prime},P_{i,1}(j+2))$ then mark $P_{i,1}(j+1)$ as skipped and set $j\coloneqq j+2$.
      2. (b)
        Else set $j\coloneqq j+1$.
  3. 3.
    Let $k$ be the maximum integer such that $P_{i,1}(k)$ is active and $k\leq j$. Mark $P_{i,1}(k)$ as passive.
  4. 4.
    Remove all skipped and passive vertices from $P_{i,1}$ to get path $P_{i,4}$.
  5. 5.
    Concatenate $P_{i,4}$ with all passive vertices after $P_{i,1}(|P_{i,4}|)$ (exclusively) on $P_{i,1}$ to get path $P_{i,5}$.
  6. 6.
    The Alter$(E)$ in Step 7 replaces each vertex $v$ on $P_{i,5}$ by $v.p$ to get path $P_{i+1}$. Set $v.p$ as active or passive according to $v$.
For any vertex $v$ on $P_{i}$ that is replaced by $v^{\prime}$ in Phase 1, if $v^{\prime}$ is not removed in Phase 4, then let $\overline{v}$ be the vertex replacing $v^{\prime}$ in Phase 6, and call $\overline{v}$ the corresponding vertex of $v$ in round $i+1$.
@@
PassivePathConstruction$(H,P)$:
Let $r=9\log b-1$. For any integer $i\in[0,r]$, let path $P_{i}$ be defined as in PathConstruction$(H,P)$.
For any integer $i\in[r,20\log b]$, given path $P_{i}$ with at least $3$ vertices at the beginning of round $i$, Expand-Maxlink$(H)$ constructs $P_{i+1}$ by the following $3$ phases:
  1. 1.
    The Alter$(E)$ in Step 2 replaces each vertex $v$ on $P_{i}$ by $v^{\prime}\coloneqq v.p$ to get path $P_{i,1}$. For any $v^{\prime}$ on $P_{i,1}$, let $\underline{v^{\prime}}$ be on $P_{i}$ such that $\underline{v^{\prime}}.p=v^{\prime}$.
  2. 2.
    After Step 6, let $v^{\prime}\coloneqq P_{i,1}(0)$, i.e., the first vertex on $P_{i,1}$. If $\underline{v^{\prime}}$ is a root at the end of round $i$ and does not increase level during round $i$ then: if the current graph contains edge $(v^{\prime},P_{i,1}(2))$ then remove $P_{i,1}(1)$. After removing all such vertices from $P_{i,1}$ we obtain $P_{i,2}$.
  3. 3.
    The Alter$(E)$ in Step 7 replaces each vertex $v$ on $P_{i,2}$ by $v.p$ to get path $P_{i+1}$.
For any vertex $v$ on $P_{i}$ that is replaced by $v^{\prime}$ in Phase 1, if $v^{\prime}$ is not removed in Phase 2, then let $\overline{v}$ be the vertex replacing $v^{\prime}$ in Phase 3, and call $\overline{v}$ the corresponding vertex of $v$ in round $i+1$.
@@
Given graph $H$ and integer $b\geq(\log n)^{100}$. For any shortest path $P$ in $H$ such that there are at most $b^{7}$ vertices on $P$, let $P_{R}$ be defined in PassivePathConstruction$(H,P)$. Let path $P_{R+1}$ be defined by replacing each vertex $v$ on $P_{R}$ with $v.p$ in Step 3 of Densify$(H,b)$, and let path $P_{\text{final}}$ be defined by replacing each vertex $v$ on $P_{R+1}$ with $v.p$ in Step 6 of Densify$(H,b)$. W.p. $1-1/(\log n)^{8}$, there are at most $2$ vertices on $P_{\text{final}}$.
@@ 5.3.1 Algorithmic Framework of Increase
Increase$(V,E,b)$:
  1. 1.
    Subgraph $H=(V,E^{\prime})=\textsc{Build}(V,E,b)$.
  2. 2.
    $(V,E_{\text{close}})=\textsc{Densify}(H,b)$.
  3. 3.
    For each vertex $v\in V$: assign a block of $b^{9}$ processors, which will be used as an indexed hash table $\mathcal{H}^{\prime}(v)$ of size $b^{9}$.
  4. 4.
    Choose a random hash function $h:[n]\to[b^{9}]$. For each vertex $v\in V$: compute $u\coloneqq v.p^{(2R+1)}$, then write $v$ into the $h(v)$-th cell of $\mathcal{H}^{\prime}(u)$ and update $v.p=u$.
  5. 5.
    For each vertex $v\in V$: if the number of vertices in $\mathcal{H}^{\prime}(v)$ is at least $2b$ then mark $v$ as a head.
  6. 6.
    For each non-loop edge $(v,w)\in E_{\text{close}}$: if $v$ is a head and $w$ is a non-head then $w.p=v$.
  7. 7.
    Shortcut$(V)$.
  8. 8.
    For each vertex $v\in V$: sample $v$ as a leader w.p. $1/2$. For each non-loop edge $(v,w)\in E_{\text{close}}$: if $v$ is a leader and $w$ is a non-leader then $w.p.p=v.p$.
  9. 9.
    Shortcut$(V)$.
  10. 10.
    Alter$(E)$.